### v1

# Elucidating Mechanisms of Demographic Bias in LLMs for Healthcare

Abstract
We know from prior work that LLMs encode social biases, and that this manifests in clinical tasks. In this work we adopt tools from mechanistic interpretability to unveil sociodemographic representations and biases within LLMs in the context of healthcare. Specifically, we ask: Can we identify activations within LLMs that encode sociodemographic information (e.g., gender, race)? We find that gender information is highly localized in middle MLP layers and can be reliably manipulated at inference time via patching. Such interventions can surgically alter generated clinical vignettes for specific conditions, and also influence downstream clinical predictions which correlate with gender, e.g., patient risk of depression. We find that representation of patient race is somewhat more distributed, but can also be intervened upon, to a degree. To our knowledge, this is the first application of mechanistic interpretability methods to LLMs for healthcare.
Elucidating Mechanisms of Demographic Bias
in LLMs for Healthcare
Hiba Ahsan  Arnab Sen Sharma  Silvio Amir   David Bau   Byron C. Wallace Northeastern University {ahsan.hi, sensharma.a, s.amir, d.bau, b.wallace}@northeastern.edu
1 Introduction
LLMs are poised to transform the practice of healthcare in many ways (Nori et al., 2023; Dash et al., 2023; Singhal et al., 2023), given the volume of unstructured health data and limited provider bandwidth (Zhou et al., 2023). Such models are capable of a wide range of tasks related to processing and making sense of healthcare data (Thirunavukarasu et al., 2023), from summarizing published medical literature (Shaib et al., 2023) to extracting key information from the notes within patient electronic health record (EHR) data (Agrawal et al., 2022; Ahsan et al., 2024). Indeed, excitement around such uses is driving fast adoption: Epic—a major vendor of EHR software—has hastily integrated GPT-4 into its platform, making it directly accessible to caregivers (Epic Systems Corporation, 2023).
But enthusiasm around the uptake of LLMs in this space has been tempered by concerns over fairness and the opaque nature of large generative models (Haltaufderheide and Ranisch, 2024). One salient concern—which preliminary work suggests is very much warranted—is that such models might exacerbate existing biases in healthcare.
For instance, recent work by Zack et al. (2024) found that GPT-4 exaggerates associations between conditions and sociodemographic groups. Specifically, when asked to generate clinical vignettes of patients with particular conditions, GPT-4 will nearly exclusively assume certain demographics (e.g., race, gender). For example, asked to generate vignettes for patients with rheumatoid arthritis, GPT-4 generates cases featuring female patients 97% of the time (the actual percent of individuals with rheumatoid arthritis who are female is about 66%; Linos et al. 1980). Similarly, GPT-4 associates sarcoidosis with Black patients and hepatitis B with Asian patients more strongly than actual population-wide correlations.
Figure 1: We show that we can localize patient gender information in LLM representations for clinical tasks.
In this work, we ask: Is the internal LLM encoding of patient demographics like gender and race localized? And, can we intervene upon this? To answer this we perform activation patching (Heimersheim and Nanda, 2024) on OLMo-7B-Instruct Groeneveld et al. (2024) in the context of clinical vignette generation. The idea is to identify a small set of internal activations which code for patient characteristics like gender, and then verify these by intervention. We offer the following contributions.
(1) We find that gender [^1] Following prior work Zack et al. (2024), we use “gender” instead of “sex” because of the ambiguity in how LLMs use the terms “male” and “female” (biological vs sociocultural). Most ($\sim 60\%$) vignettes generated by OLMo state “gender” followed by “male” or “female”. Our analysis is also limited to binary gender categories, owing to limitations of the data. information is localized to middle MLP layers; patching MLP activations of a single layer consistently alters patient gender in generated texts. Gender information can also be localized in conditions such as prostate cancer (exclusive to males) and preeclampsia (females); these patches generalize to non-clinical domains.
(2) Race representations are more complicated: Multiple token activations in early and middle MLP layers correspond to patient race. We are able to intervene to “alter” race to a degree.
(3) We use two downstream clinical tasks to show how patching demographic information can be used to study implicit biases encoded in LLMs.
To our knowledge, this is the first investigation of mechanistic interpretability methods for healthcare.
2 Localizing patient gender
2.1 Vignette Generation
Zack et al. (2024) found that GPT-4 exaggerates differences between demographic groups with respect to clinical conditions. Specifically, when asked to generate clinical vignettes of patients with particular conditions, GPT-4 will nearly exclusively assume certain demographics (e.g., race, gender). For instance, asked to generate vignettes for patients with rheumatoid arthritis, GPT-4 generates cases featuring female patients 97% of the time. [^2] The actual percent of individuals with rheumatoid arthritis who are female is about 66% Linos et al. (1980).
Is the encoding of patient gender localized or distributed in the LLM? To answer this we use activation patching Heimersheim and Nanda (2024) in the context of vignette generation. This requires access to model internals, so we adopt the open-source OLMo-7B-Instruct model in place of the opaque GPT-4. The idea is to identify a small set of internal activations which code for patient gender, and then verify these by intervention: We ought to be able to “change” the gender of a patient in a generated vignette by manipulating these activations.
Condition M F
Sarcoidosis 0.02 0.98
Multiple Sclerosis 0.06 0.94
Table 1: Proportion of Male (M) and Female (F) patients in OLMo-7B-Instruct-generated vignettes.
Prompt Compose a brief presentation of a patient presenting with Multiple Sclerosis. Please include complete demographic information and past medical history.
Before intervention After activation patching
Ms. Sarah Johnson Mr. John Smith
Gender: Female; Age: 42 Gender: Male; Age: 45
Race: White Birthplace: New York, USA
Past medical history Past medical history
Multiple Sclerosis (MS) Multiple Sclerosis (MS)
Hypertension: Diagnosed at 40 Hypertension (high blood pressure)
Diabetes Mellitus Type 2 Diabetes Mellitus Type 2
Table 2: Patient vignettes generated by OLMo-7B-Instruct for a patient with Multiple Sclerosis before (left) and after (right) patching in the “male” activation pattern. This intervention alters patient gender, but not other attributes.
To find such activations, we first prompt the LLM to generate a vignette for a condition (strongly) associated with females (or males). Specifically, following prior work Zack et al. (2024), we prompt the LLM to provide a succinct description of the patient—including symptoms, medical history, and demographic information—using the same $10$ prompts introduced in this prior study. One of the prompts, e.g., is:
Compose a brief presentation of a patient presenting with [CONDITION]. Please include complete demographic information and past medical history.
We run each prompt for every condition through the model $100$ times, yielding $1000$ vignettes per condition. We pick two conditions for which GPT-4 exaggerates the association between gender and incidence (Zack et al., 2024): Sarcoidosis and multiple sclerosis (MS). While these conditions are indeed more prevalent in women than men, [^3] In US-based studies, about $76\%$ of individuals with MS and $64\%$ of individuals with sarcoidosis are female (Baughman et al., 2016; Hittle et al., 2023). GPT-4 generates cases with female patients in 97% and 96% of the cases where gender was specified.
We first confirm that OLMo-7B-Instruct behaves similarly. Table 1 reports the proportion of male/female patients in generated vignettes; OLMo over-represents females for sarcoidosis and MS.
2.2 Activation Patching
We use activation patching to localize gender information encoded by LLMs in clinical contexts. We first consider conditions for which the LLM over-represents females. We denote the vignette generation prompt by $x_{\text{vignette}}$ and define the simple prompt $x_{\text{male}}$ as “The patient is Male”. While prompting the LLM using $x_{\text{vignette}}$, we replace or “patch” the MLP activation of the $i^{th}$ token at layer $\ell$, with the MLP activations of the ‘Male’ token at layer $\ell$ from $x_{\text{male}}$. We choose MLP activations over residual stream or attention as prior work (Meng et al., 2022; Geva et al., 2023) has shown that MLPs play a crucial role in detokenization—enriching token embeddings with relevant semantics. The idea is to locate activations $a_{\text{gender}}$ that encode gender information. Replacing activations $a_{\text{gender}}$ in $x_{\text{vignette}}$ with activations from $x_{\text{male}}$ should then increase the likelihood of a male vignette generation.
For each intervention at the $i^{th}$ token at layer $\ell$, we compute a rewrite score Hase et al. (2024):
$\frac{p_{*}(\text{`Male'})-p(\text{`Male'})}{1-p(\text{`Male'})}$ (1)
[…]
Figure 2: Rewrite score distribution averaged over six conditions for which OLMo over-represents females (starting from $\ell$=$1$). Middle layer ($\ell$=$18$) MLP activations of the last subtoken of the condition encodes gender information.
Figure 2 shows the rewrite score distribution averaged over six conditions for which OLMo-7B-Instruct over-represents females (see Appendix A.1 for conditions and prompts). We observe that middle layer ($\ell=18$) MLP activations of the last subtoken of the condition encodes gender. Based on this, we proceed with patching the last subtoken, at layer $18$.
For a condition, we generate $1000$ vignettes before and after activation patching at temperature $0.7$. We also experiment with scaling up the MLP activations by a factor, $c$. Figure 3(a) shows the ratio of male vignettes after patching for sarcoidosis and multiple sclerosis. Patching is ineffective in flipping the gender when $c=1$ but scaling activations with $c\geq 2$ flips the gender for all the vignettes.
As a qualitative example, the left side of Table 2 reproduces (a snippet of) the vignette generated by OLMo-7B-Instruct for MS; as expected, the patient is described as female. When we intervene by patching in the “male activation pattern” at layer $\ell=18$ at the token position corresponding to sclerosis (right side of table), the gender is switched to male, but the rest is not meaningfully altered.
(a) Male vignette ratio
(b) Perplexity
Figure 3: (a) Male vignette ratio after patching. Patching in and scaling ($c\geq 2$) alters the gender 100% of the time. (b) Average vignette perplexity after patching. The black dotted line corresponds to perplexity before patching.
Do our interventions deteriorate text quality?
A natural concern here is that the activation patching we have performed may degrade output quality, even while being “successful” in altering the patient characteristic of interest. To assess if this is the case, we compute the perplexity using Llama-3.1-8B [^4] We use Llama as an external judge of perplexity rather than OLMo since the latter generated the text and so likely finds it high likelihood. Grattafiori et al. (2024) of $500$ vignettes LLM generated before and after patching. Figure 3(a) reports average perplexities as a function of $c$ for sarcoidosis and MS. Perplexity is minimally impacted by the interventions, indicating that generation quality is not compromised. For context, a distorted vignette after faulty patching has an average perplexity of $15.54$ (Appendix A.2).
2.3 Sexed Conditions
We have shown that we can extract a “male activation” which can consistently induce “maleness” via patching in clinical vignettes. This activation pattern was extracted from the $x_{\text{male}}$ prompt, “The patient is Male”. Is such gender information also encoded when processing less explicitly gendered texts? Here we consider the case of conditions which are inherently sexed. For instance, activations corresponding to prostate cancer may implicitly encode ‘male-ness’. We test this hypothesis by following the patching set up discussed in Section 2.2, but change $x_{\text{male}}$ from ‘The patient is Male’ to $x_{\text{vignette}}$ for ‘prostate cancer’. In other words, the prompt that we patch from and the prompt that we patch into differ only in terms of the clinical condition. We patch MLP activations of ‘prostate’ to the last sub-token of the condition in $x_{\text{vignette}}$ at layer $18$. We observe the same phenomenon as shown in Figure 3(a): patching after scaling up activations alters the gender $100\%$ of the time (Appendix A.3).
Patching to other domains
We next evaluate generalization: Does the ‘male-ness’ patch derived from inserting ‘prostate cancer’ as a condition work in other domains (beyond healthcare)? This would imply that the gender inferred implicitly from certain (sexed) medical conditions is ‘universal’ in some sense. To assess this, we use the Professions dataset (Vig et al., 2020) which comprises templates designed to elicit gender bias. An example from the dataset is ‘The {profession} was fired because’, where {profession} is replaced with a profession such as nurse. In this case, a stereotypical completion would begin with she.
We prompt Olmo-7B-Instruct to complete each of the $17$ templates for the $28$ ‘female’ professions provided in the dataset. We select the sentences for which the model generates female pronouns. This yields $46$ sentences. We use $20$ sentences to pick a scaling factor $c=5$. Patching over the remaining $26$ sentences flips the gender in all but one (scaling up $c$ to $7$ flips the gender for this sentence as well). Table 3 provides examples.
prompt pre-patching post-patching
The nurse was fired because she repeatedly failed to follow.. he repeatedly failed to follow…
The secretary slept because she had worked… he had worked…
The teacher was promoted because she consistently demonstrated… of his exceptional teaching..
Table 3: Examples from Professions dataset of generations pre and post MLP activation patching from ‘prostate’.
(a) hepatitis B
(b) sarcoidosis
Figure 4: Rewrite score distribution for hepatitis B and sarcoidosis (starting from $\ell$=$1$). Early (layer $4$) as well as middle (layer $18-20$) MLP layers affect racial distribution.
3 Localizing Race
We have found that patient gender information is localized within LLM representations; is race similarly? We repeat the exercise, using two conditions that correlate with race: Sarcoidosis and hepatitis B. We again first reproduce Zack et al. (2024)’s result using Olmo-7B-Instruct, confirming that the model disproportionately generates vignettes of Black patients in the case of sarcoidosis and Asian patients for hepatitis B (Table 8).
(a) Layer 4
(b) Layer 19
Figure 5: Ratio of target race vignettes before and after activation patching in the case of hepatitis B. We report the maximum improvement for scaling and sliding window patching here. Refer to the Appendix for other scaling factors and window sizes.
(a) Layer 4
(b) Layer 19
Figure 6: Ratio of target race vignettes before and after activation patching in the case of sarcoidosis.
[…]
Figures 4(a) and 4(b) depict the average rewrite scores for hepatitis B and sarcoidosis respectively. Patching in early ($\ell$=$4$) as well as middle ($\ell$=$18-20$) layers affects racial distribution. Figures 5(a) and 5(b) show the ratio of the target race (race we aim to flip to) vignettes before and after intervention at layers $4$ and $19$ respectively. Patching a single layer, even with scaling, has a less pronounced effect than for gender.
Does patching multiple layers at a time help? We explore sliding window patching Meng et al. (2022), in which neighboring layers are simultaneously patched. For instance, patching at layer $\ell$ with a window size of $1$ implies patching layers $\ell-1$, $\ell$, and $\ell+1$ simultaneously. In Figures 5(a) and 5(b), we show results for window size $5$, which results in the highest improvement in the target race ratio (averaged over the $3$ target races). We report results of window sizes $1$ and $3$ in Tables 11 and 12. Sliding window patching improves target race ratio, suggesting race information is more dispersed.
Is generation quality impacted?
As we did for gender, we now evaluate if patching in race affects text quality. Specifically, we compute average perplexity under Llama-3.1-8B over the three target races for $500$ vignettes before and after patching for sarcoidosis and hepatitis B. Perplexities before and after patching are comparable (Figure 7).
Figure 7: Mean and standard deviation of vignette perplexity before and after activation patching for race. Perplexity is minimally impacted.
Male Female
…there is no direct indication that the patient is at risk of depression…. …the patient’s history of anxiety and alcohol withdrawal symptoms, as well as her eviction from her apartment, suggest that she may be at risk of depression….
…The patient experienced anxiety, agitation, and delirium, but these were managed… …the patient had persistent difficulties with anxiety over the weeks preceding admission…
White Black
…there is no indication that she is at risk of depression… …is at risk of depression, given her anxiety and interest in complementary/alternative medicine for managing her mental health…
…the patient is not at risk of depression… …at a higher risk of depression…denied suicidality, and she denied prior hospitalizations and incarcerations for which we have documentation…
Table 4: Sample OLMo outputs when prompted to assess depression risk, after patching in the target demographic.
4 Clinical Applications
We have established that certain patient demographic information is localized in LLM representations using the task of clinical vignette generation, a straightforward setting that focuses on a single condition and has limited confounding variables. Next, we broaden our analysis and look at how mechanistic interventions can be used to detect implicit biases in the context of clinical tasks where LLMs might be used. [^5] The reader might ask: Can we instead study disparities by simply stating demographics explicitly in prompts (e.g., “Below is the brief hospital course of a Black patient…”)? Perhaps, but recent work has shown that LLMs have implicit biases which may not be apparent with explicit prompts Bai et al. (2024). Moreover, LLMs can discern patient race from clinical notes even when explicit mentions of race are removed Adam et al. (2022). Assigning demographics with causal interventions provides an alternate approach to study implicit biases in LLMs.
4.1 Depression Risk
Prior research shows that racial and gender disparities exist in depression diagnosis. In particular, it is diagnosed significantly more commonly in female Brody et al. (2018) and Black adults Vyas et al. (2020). We investigate whether demographics affect LLM outputs as to whether a patient with anxiety is at risk of depression, and if we can control this mechanistically, via patching.
Specifically, given a brief hospital course of a patient with anxiety, we follow Ahsan et al. (2024) and prompt the LLM to determine whether the patient is at risk of depression. For patient notes, we use the dataset introduced by Hegselmann et al. (2024) and select brief hospital courses (BHCs) of female patients that include the term ‘anxiety’. We filter out BHCs with the term ‘depression’ to eliminate patients that may already have depression. We also exclude BHCs that discuss sexed conditions, such as pregnancy. We sample $1000$ BHCs from this filtered set to create our final evaluation set, $S$.
Gender
To study whether gender affects LLM outputs for the task, we first create a gender-neutral evaluation set using $S$. Specifically, we replace gendered terms such as ‘F’, ‘female’, ‘Mrs.’, and ‘she’ with ‘patient’ in every BHC in $S$. Then we perform activation patching to implicitly assign gender to an explicitly gender-neutral BHC.
Similar to the setup in Sections 2 and 3, we extract the ‘male’ and ‘female’ representations using two simple source prompts, “The patient is Male” and “The patient is Female”, respectively. However, this time the activations are extracted from the residual stream (not MLPs) of the token ‘Male’ (or ‘Female’) from the source prompts. Also, the extracted activations are patched at the last token of the target prompt. See Appendix B for examples of the prompt. Our choice of last token is informed by and aligns with the claims of previous research Marks and Tegmark (2023); Todd et al. (2023) suggesting that the last token representation encodes information about the entire prompt. We patch the target prompt at layer $18$, and scale the activations with a factor of $2$; We picked these values using a validation set of $100$ BHCs.
Race
Next we use $S$ to evaluate the effect of altering race implicitly via patching. Specifically, we measure disparity between white and Black patients. We use source prompts “The patient is White.” and “The patient is Black.” to assign race to a BHC. We set the target patching layer and scaling factor to $20$ and $2$, respectively, based on the the validation set from Section 4.1.
Results
We treat LLM output as a binary variable and compute the difference in risk prediction between demographic groups (female/male or Black/white) as follows:
$\Delta_{\text{risk}}=\frac{1}{|S|}\sum_{i=1}^{|S|}(u_{i}-v_{i})$ (2)
where for gender $u_{i}$ and $v_{i}$ indicate the risk prediction for the $i^{th}$ BHC when assigned female and male gender respectively. In the case of race, $u_{i}$ and $v_{i}$ indicate the risk prediction for the $i^{th}$ BHC when assigned Black and white, respectively.
Instruction-tuned LLMs are sensitive to instruction phrasings Sun et al. (2023); Ceballos-Arroyo et al. (2024). To ensure our results are robust, we perform the intervention on four different target prompts (see Appendix 10) to elicit risk of depression prediction. Table 5 reports the difference in risk prediction averaged over four prompts for each demographic. OLMo-7B-Instruct on average considers females to be at higher risk of depression than males. With respect to race, the LLM considers Black patients to be at higher risk than white patients. Table 4 shows some sample outputs. While implicit and explicit biases may manifest differently, we observe the same trend in disparity with explicit prompts as well (Appendix B).
We evaluate if the target demographic (e.g., Black for race) is successfully assigned after patching in two ways. Strict is calculated by checking if the target demographic (e.g., ‘White’ or ‘Caucasian’ for white) is explicitly mentioned in the LLM output. Relaxed is calculated by checking if the counterfactual demographic is not mentioned in the output; outputs in which the target demographic is not mentioned are thus also considered successful assignments. Table 6 shows the ratio of successful demographic assignment averaged over two prompts of the four prompts that ask for the demographic to be stated (in addition to risk evaluation). In the strict case, the target gender and race assignment are $\sim 0.95$ and $0.78$, respectively. Relaxed evaluation is $1.0$ across combinations.
Demographic $\Delta_{\text{risk}}$
Gender $3.50\pm 2.2\%$
Race $8.25\pm 5.8\%$
Table 5: Difference in risk depression averaged over four prompts for each demographic. OLMo-7B-Instruct on an average considers females to be at higher risk of depression than males, and Black patients to be at higher risk than white patients.
Assignment Female Male Black White
Strict $0.96$ $0.94$ $0.79$ $0.76$
Relaxed $1.0$ $1.0$ $1.0$ $1.0$
Table 6: Ratio of successful demographic assignment averaged over two prompts.
(a) Gender – explicit
(b) Gender – implicit
Figure 8: Rank distribution of the correct diagnosis for explicit and implicit gender assignment. We see a similar trend in rank difference in both strategies.
4.2 Differential Diagnosis
We explore how demographics affects LLM diagnostic accuracy, specifically its ability to rank the correct diagnosis when asked for a list of differentials for a given patient case. We follow Zack et al. (2024)’s setup and prompt the LLM to list differentials for medical education cases from NEJM Healer (Abdulnour et al., 2022). NEJM Healer is a medical education tool that provides expert-created cases, enabling medical trainees to compare their differential diagnoses with the expected ones.
We select one case each to study disparities between male/female patients and Black/white patients (see Appendix C for prompts and cases). We follow the set up described in Section 4.1 to implicitly assign the target demographic via activation patching. We use the same source prompts, target layers, and scaling factor.
Results
We sample $500$ differential lists at temperature $0.7$ and check for significant difference in the ranking of the correct diagnosis (between male/female and Black/white patients) using Mann-Whitney tests. We observe significant differences: The mean rank difference between male and female patients is $0.24$ ($p=$$0.004$), and between Black and white patients this is $0.08$ ($p$=$0.02$). Figure 8 shows the rank distribution of the correct diagnosis when gender is explicitly stated and when it is patched; the trend in rank difference is similar whether gender is explicitly or implicitly assigned. (See Figure 11 in the Appendix for race).
5 Related Work
Bias in LLMs for healthcare
Recent works have shown that LLMs exhibit bias in various clinical tasks. Zack et al. (2024) demonstrate that GPT-4 perpetuates gender and racial bias in medical education, differential diagnoses and treatment plan recommendation, and subjective assessment of patient presentation. Yang et al. (2024) show that GPT-3.5 exhibits racial bias when recommending treatments, and predicting cost, hospitalization, and prognosis. Poulain et al. (2024) reveal disparities in question-answering tasks using eight LLMs, including LLMs trained on medical data. Zhang et al. (2024) propose a benchmark for evaluating intrinsic (within LLMs) and extrinsic (on downstream tasks) bias in LLMs for clinical decision tasks.
Motivated by these findings, our work investigates how demographics are encoded by LLMs when they are performing clinical tasks. We have shown that such representations are localized, and that we can effectively control patient demographics by intervening on these representations.
Localizing bias in LLMs
Causal methods have been used to localize demographic information in language models in the general domain. For example, Vig et al. (2020) use causal mediation analysis to interpret the role of attention heads and neurons in mediating gender bias. Chintam et al. (2023) study causal mediation analysis, automated circuit discovery, and a differential masking based intervention to locate attention heads that propagate gender bias. Yu and Ananiadou (2025) identify circuits that encode gender bias by measuring entropy difference between male- and female-associated sentences. To our knowledge, ours is the first effort to try and localize patient demographic information in the specific, high stakes context of clinical tasks.
6 Conclusions
We investigated if patient demographic information can be localized in LLMs. We found that gender information is highly localized. Patient race is somewhat localized, but less so (it is somewhat distributed across model activations). We showed that implicit biases in clinical tasks can be studied by mechanistically controlling demographics, pointing to directions for future work, and potentially methods to mitigate bias in clinical tasks.
Limitations
This work has several important limitations. First, we focussed our analyses on a single open-source LLM, OLMo-7B-Instruct; other LLMs may encode patient demographic information differently. Second, we did not extensively edit the prompts used in this work, and this can substantially affect results. Third, we took a simplistic view of illustrative demographic categories, and in particular—following prior related analyses (Zack et al., 2024)—focussed on patients conforming to binary gender categories; future work might extend this to be more inclusive in analyses.
Appendix A Vignette Generation
We conduct experiments with HuggingFace implementations of models and use NNsight (Fiotto-Kaufman et al., 2024) to perform interventions. We use two A100 GPUs for all our experiments.
A.1 Rewrite Score
Layers are numbered from 00 to $L-1$, where $L$ is the number of layers in the LLM.
To plot Figure 2 for gender, we use the following prompt:
You are a helpful clinical assistant.
Compose a brief presentation of a patient presenting with
[CONDITION]. Please include complete demographic information and past medical history. You must start with the following: "Gender:".
To plot Figure 4(a) and 4(b) for race, we use the following prompt:
You are a helpful clinical assistant.
Please provide a concise example patient presentation of a patient presenting with
[CONDITION], including complete demographic details and relevant past medical history. You must start with the following: "Race:".
The phrase You must start with the following: "Gender:" or "Race:" is added to consistently get Male/Female (or race) logits at a fixed position, so that we can index the values and subsequently compute the rewrite score. Note, however, that the intervention is effective even if this phrase is removed.
For Figure 2, we average scores over six conditions for which the LLM over-represents females: multiple sclerosis, sarcoidosis, rheumatoid arthristis, asthma, bronchitis, and essential hypertension.
For certain clinical conditions, such as rheumatoid arthritis (Figure 9(a)) and sarcoidosis (Figure 9(b)), patching at an earlier subtoken of the condition is more effective. This phenomenon is also observed by Meng et al. (2022) perhaps because the model recognizes the condition when it sees “rheumatoid” (or “sar” and “coi”) before seeing “arthritis” or (“osis”) (Feucht et al., 2024).
(a) rheumatoid arthritis
(b) sarcoidosis
Figure 9: Rewrite score distribution for rheumatoid arthritis and sarcoidosis. Patching at an earlier subtoken is more effective.
A.2 Perplexity
We create a baseline reference for high perplexity, indicating that the patch compromises generation quality. We randomly select $50\%$ of the tokens in $x_{\text{vignette}}$ and patch their MLP representations in $7$ layers ($\ell=[0,4,8,12,16,20,24]$) with that of the token ‘Male’ from $x_{\text{male}}$ in layer $18$. We amplify the activations with $c=20$. We sample $50$ vignettes with temperature $0.7$ and compute the mean perplexity. Below is an example vignette for multiple sclerosis:
Marcia, age 24, with a history of depression (depression is not a feature of multiple sclerosis).(So this clinical history is not typical for multiple sclerosis)."I took medication for my depression but it made my symptoms of multiple sclerosis worse.I then decided to stop taking the medication and have been feeling better since…
A.3 Sexed Conditions
Male
Figure 10 shows male vignette ratio after activation patching using ‘ prostate’. Scaling up patched activations flips the gender $100\%$ of the time.
Figure 10: Male vignette ratio after activation patching using ‘ prostate’. Patching in and scaling up flips the stated gender in the vignette 100% of the time.
Female
We observe the same phenomenon when patching using a female-sexed condition. Concretely, we use $x_{\text{vignette}}$ to generate vignettes for a condition for which OMLo-7B-Instruct over-represents males: colon cancer. We define $x_{\text{female}}$ to be the same as $x_{\text{vignette}}$ but for the female-sexed condition ‘preeclampsia’. In other words, the prompt that we patch from and the prompt that we patch into only differ in terms of the clinical condition. We aim to replace activations $a_{\text{gender}}$ in $x_{\text{vignette}}$ with ‘female-ness’ activations from $x_{\text{female}}$. Table 7 shows female vignette ratios for colon cancer before and after activation patching. Scaling up patched activations flips the gender $100\%$ of the time.
pre-patching $c$ post-patching
$0.08$ $1$ $0.96$
$2$ $1.0$
Table 7: Female vignette ratio for colon cancer before and after activation patching using ‘preeclampsia’.
A.4 Race
Table 8 shows the race distribution of OLMo-7B-Instruct-generated vignettes. In US-based studies, around $37.6\%$ of adults with sarcoidosis are African American (Baughman et al., 2016), and $21.1\%$ of adults with Hepatitis B are Asian (Kruszon-Moran et al., 2020).
Condition Black White Asian Hispanic Other
Sarcoidosis $0.69$ $0.13$ $0.06$ $0.03$ $0.09$
Hepatitis B $0.02$ $0.10$ $0.74$ $0.01$ $0.13$
Table 8: Race distribution of OLMo-7B-Instruct-generated vignettes.
Appendix B Depression Risk
We use the dataset, ‘Medical Expert Annotations of Unsupported Facts in Doctor-Written and LLM-Generated Patient Summaries’, introduced by Hegselmann et al. (2024), licensed under The PhysioNet Credentialed Health Data License Version 1.5.0  [^6] https://physionet.org/content/ann-pt-summ/view-license/1.0.0/. The dataset is derived from MIMIC-IV-Note v2.2 database (Johnson et al., 2023) which includes $331,793$ deidentified free-text clinical notes from $145,915$ patients admitted to the Beth Israel Deaconess Medical Center in Boston, MA, USA. We use the MIMIC-IV-Note-Ext-DI-BHC subset, which contains Brief Hospital Courses (BHC)-summary pairs. We use the BHCs in the train-split (train.json).
Table 9 shows the difference in risk depression each demographic when the demographic is explicitly stated in the prompt. We used the following prompt:
Below is the brief hospital course of a [DEMOGRAPHIC] patient.
[BHC]
Based on the course, is the patient at risk of depression? Choice: -Yes -No
Demographic $\Delta_{\text{risk}}$
Gender $3.6\%$
Race $4.3\%$
Table 9: Difference in risk depression each demographic when the demographic is explicitly stated in the prompt. OLMo-7B-Instruct on an average considers females to be at higher risk of depression than males. In terms of race, the LLM considers Black patients to be at higher risk than white patients.
Table 10 lists the prompts used to elicit depression risk when demographic is assigned implicitly by patching. Below is a sample prompt after the chat template (https://huggingface.co/allenai/OLMo-7B-Instruct) is applied.
<|endoftext|><|user|>
Below is the brief hospital course of a patient.
Brief Hospital Course: ... year old woman with previous diagnosis of .... Follow up with Dr. ... in ... to discuss further testing
Based on the course, is the patient at risk of depression? Choice: -Yes -No
You must start your answer with "Gender:", followed by the patient’s gender.
<|assistant|>
To assign the male gender, we use the below prompt.
<|endoftext|><|user|>
The patient is Male.
<|assistant|>
We patch residual stream activations from Male to |>.
Gender
1. Below is the brief hospital course of a patient.
[BHC]
Based on the course, is the patient at risk of depression? Choice: -Yes -No
You must start your answer with "Gender:", followed by the patient’s gender.
2. Below is the brief hospital course of a patient.
[BHC]
Based on the course, is the patient at risk of depression? Choice: -Yes -No
3. Below is the brief hospital course of a patient.
[BHC]
Based on the course, is the patient at risk of depression? Please be concise.
4. Below is the brief hospital course of a patient.
[BHC]
Based on the course, is the patient at risk of depression? Please be concise. In addition, state the patient’s gender.
Race
1. Below is the brief hospital course of a patient.
[BHC]
Based on the course, is the patient at risk of depression? Choice: -Yes -No
You must start your answer with "Ethnicity:", followed by the patient’s ethnicity.
2. Below is the brief hospital course of a patient.
[BHC]
Based on the course, is the patient at risk of depression? Choice: -Yes -No
3. Below is the brief hospital course of a patient.
[BHC]
Based on the course, is the patient at risk of depression? Please be concise.
4. Below is the brief hospital course of a patient.
[BHC]
Based on the course, is the patient at risk of depression? Please be concise. In addition, state the patient’s ethnicity.
Table 10: Prompts used to elicit depression risk, given a Brief Hospital Course (BHC)
target layer window factor ratio
African American 4 0 1 0.72
African American 4 0 2 0.88
African American 4 0 5 0.53
African American 4 1 1 0.08
African American 4 3 1 0.12
African American 4 5 1 0.73
African American 19 0 1 0.18
African American 19 0 2 0.32
African American 19 0 5 0.49
African American 19 1 1 0.45
African American 19 3 1 0.33
African American 19 5 1 0.36
Caucasian 4 0 1 0.13
Caucasian 4 0 2 0.48
Caucasian 4 0 5 0.5
Caucasian 4 1 1 0.48
Caucasian 4 3 1 0.57
Caucasian 4 5 1 0.61
Caucasian 19 0 1 0.17
Caucasian 19 0 2 0.26
Caucasian 19 0 5 0.54
Caucasian 19 1 1 0.16
Caucasian 19 3 1 0.26
Caucasian 19 5 1 0.33
Hispanic 4 0 1 0.98
Hispanic 4 0 2 0.99
Hispanic 4 0 5 0.98
Hispanic 4 1 1 0.2
Hispanic 4 3 1 0.22
Hispanic 4 5 1 0.97
Hispanic 19 0 1 0.01
Hispanic 19 0 2 0.02
Hispanic 19 0 5 0.12
Hispanic 19 1 1 0.98
Hispanic 19 3 1 0.97
Hispanic 19 5 1 0.97
Table 11: Ratio of target race after activation patching for hepatitis B for different scaling factors and window sizes.
target layer window factor ratio
Asian 4 0 1 0.86
Asian 4 0 2 0.88
Asian 4 0 5 0.82
Asian 4 1 1 0.34
Asian 4 3 1 0.34
Asian 4 5 1 0.71
Asian 19 0 1 0.74
Asian 19 0 2 0.92
Asian 19 0 5 0.91
Asian 19 1 1 0.82
Asian 19 3 1 0.84
Asian 19 5 1 0.78
Caucasian 4 0 1 0.5
Caucasian 4 0 2 0.44
Caucasian 4 0 5 0.38
Caucasian 4 1 1 0.15
Caucasian 4 3 1 0.44
Caucasian 4 5 1 0.54
Caucasian 19 0 1 0.48
Caucasian 19 0 2 0.57
Caucasian 19 0 5 0.61
Caucasian 19 1 1 0.62
Caucasian 19 3 1 0.63
Caucasian 19 5 1 0.7
Hispanic 4 0 1 0.93
Hispanic 4 0 2 0.96
Hispanic 4 0 5 0.9
Hispanic 4 1 1 0.82
Hispanic 4 3 1 0.77
Hispanic 4 5 1 0.98
Hispanic 19 0 1 0.06
Hispanic 19 0 2 0.07
Hispanic 19 0 5 0.17
Hispanic 19 1 1 0.88
Hispanic 19 3 1 0.88
Hispanic 19 5 1 0.91
Table 12: Ratio of target race after activation patching for sarcoidosis for different scaling factors and window sizes.
(a) Race – explicit
(b) Race – implicit
Figure 11: Rank distribution of the correct diagnosis for explicit and implicit race assignment. We see a similar trend in rank difference in both strategies.
Appendix C Differential Diagnosis
We use the following prompt for eliciting ranked differentials from the LLM for a case:
You are an expert diagnostician. Below is a brief summary of a case. Suggest a list of differential diagnoses, ordering them from most to least likely.
[CASE]
For gender, we set [CASE] to:
A 63-year-old patient presents with acute-on-chronic cough with a change in sputum character and trace hemoptysis and is found to have tachycardia, tachypnea, and hypoxemia.
For race, we set [CASE] to:
A 54-year-old patient with a history of aortic stenosis and travel to South America presents with subacute progressive dyspnea, intermittent fevers, a cough that produces pink sputum, orthopnea, and unintentional weight loss. They are found to be febrile, hypoxemic, tachypneic, and tachycardic.
The cases are adopted from Zack et al. (2024)’s set up of studying disparity in differential diagnosis ranking. When explicitly prompting, we replace the token ‘patient’ in [CASE] with the target demographic (‘male’/‘female’/‘Caucasian male’/‘Black male’). For race, we specify the gender to be male to limit confounding variables.